can someone plss help me by monday? im really stuck :(
solve the equation.
13. x-2/x+2 = x+10/x

Answer by stanbon(75887)

x-2/x+2 = x+10/x
---
Cross-multiply:
(x-2)x = (x+2)(x+10)
-------
x^2-2x = x^2+12x+20
----
-2x = x^2+12x+20
---
x^2+14x+20 = 0
---
Use Quadratic Formula:
x = [-14 +- sqrt(14^2-4*1*20)]/2
----
x = [-14 +- sqrt(116)]/2
----
x = [-14 +- 2sqrt(29)]/2
==============================
x = [-7+sqrt(29)] or x = [-7-sqrt(29)]
=====================
Cheers,
Stan H.
================